# Efficient Redundancy Reduction for Open-Vocabulary Semantic Segmentation

### v1

@@ III-C Channel Reduction Module
The pipeline of the function $\Phi$ is illustrated in Fig. 4. Initially, $S$ is averaged across the $C$ dimension, resulting in $D\in\mathbb{R}^{K\times H_{c}\times W_{c}}$, representing the probability of each category at each spatial location. Subsequently, for each spatial location, the top $k$ classes are selected based on their probability values and arranged in descending order:
$N=Top(D,k),$ (5)
@@
where $N\in\mathbb{R}^{k\times H_{c}\times W_{c}}$ represents the indices of the top $k$ categories. Then we calculate the scores for each category. The score for $t$-th category, denoted as $R_{t}$, is computed as:
$R_{t}=\sum_{c=1}^{k}{\lambda^{c-1}\left(\sum_{i=1}^{H_{c}}{\sum_{j=1}^{W_{c}}{\left[N_{c}(i,j)=t\right]}}\right)},$ (6)
@@
where $\lambda^{c-1}$ represents the score assigned to the class ranked $c$. $\lambda$ is the score coefficient and is expected to meet the condition $10<\lambda<$ 1, as higher-ranked categories should receive higher scores. The notation []delimited-[][$\cdot$][ ⋅ ] represents a logical expression that yields 1 when the condition is true and 0 otherwise. After scoring all categories, we obtain =$\mathcal{R}=\left\{R_{i}\right\}_{i=1}^{K}$. Finally, $\mathcal{R}$ is sorted in descending order, and only the top $P$ categories are retained. Subsequently, $S_{r}$ is utilized instead of $S$ in the following calculations.
@@
The original cost maps $S$ comprise the existing and redundant classes. Let $\mathcal{Q}_{e}=\{S(i)\}_{i=1}^{M}$ collect the original cost maps corresponding to the existing classes, where $M$ is the number of existing classes in $I$. The cost maps corresponding to the redundant classes are represented by $\mathcal{Q}_{r}=\{S(i)\}_{i=M+1}^{K}$, and those corresponding to the redundant classes after partial reduction are denoted as $\mathcal{Q}_{p}=\{S(i)\}_{i=M+1}^{P}$. In the subsequent subsections, we will analyze the impact of $\mathcal{Q}_{r}$ and $\mathcal{Q}_{p}$ on class-level and spatial-level attention mechanisms (the primary operations after CRM), as well as on latency.
III-C1 Impact on class-level attention.
The class-level attention is the attention mechanism applied to the class dimension of $S$, which is used in CFM as visualized in Fig. 3(b). Its output of the $q$-th class in $\mathcal{Q}_{e}$ before reduction is:
$O^{e,r}(q)=\sum_{k=1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)},$ (7)
@@
$O^{r}(q)=\sum_{k=M+1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)}.$ (8)
@@
$\Delta_{p}=\frac{O^{p}(q)}{O^{e,p}(q)}<\Delta_{r}=\frac{O^{r}(q)}{O^{e,r}(q)}$ (9)
Proof.
$\Delta_{r}=\sum_{k=M+1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}.$ (10)
Similarly, $\Delta_{p}$ can be simplified as:
$\Delta_{p}=\sum_{k=M+1}^{P}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{P}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)},$ (11)
$\Delta_{r}^{\prime}=\left(\frac{1}{\Delta_{r}}-1\right),$ (12)
$\Delta_{p}^{\prime}=\left(\frac{1}{\Delta_{p}}-1\right).$ (13)
Since $10<\Delta_{r},\Delta_{p}<$ 1, it follows that $\Delta_{r}^{\prime},\Delta_{p}^{\prime}>0$. To compare $\Delta_{r}$ and $\Delta_{p}$, we first compare $\Delta_{r}^{\prime}$ and $\Delta_{p}^{\prime}$, by analyzing the following formula:
$\frac{\Delta_{r}^{\prime}}{\Delta_{p}^{\prime}}=\frac{\sum_{k=1}^{M}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}}{\sum_{k=M+1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}}\frac{\sum_{k=M+1}^{P}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}}{\sum_{k=1}^{M}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}}.$ (14)
Let us define:
$\alpha=\sum_{k=1}^{M}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{M}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}$ (15)
and
$\beta=\sum_{k=M+1}^{P}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=M+1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}.$ (16)
Thus, we have
$\frac{\Delta_{r}^{\prime}}{\Delta_{p}^{\prime}}=\alpha\beta.$ (17)
$\forall 1\leq i,j\leq P,A_{ij}=B_{ij}.$ (18)
This implies that $\alpha=1$ and $\beta<1$. Hence,
$\frac{\Delta_{r}^{\prime}}{\Delta_{p}^{\prime}}=\alpha\beta<1.$ (19)
@@
The above inequality suggests that as the number of redundant classes increases, they contribute more to the final output. Thus, existing classes contribute less. As such, we can deduce that reducing redundancy facilitates the class-level attention to focus more on existing classes.
Figure 5: Architecture of our proposed SFM and CFM. SSR signifies Spatial-level Sequence Reduction, while CSR represents Class-level Sequence Reduction. DW-Conv refers to deep-wise convolution, and FC indicates fully connected layers.
III-C2 Impact on spatial-level attention
@@
Spatial-level attention aims to capture dependencies among different pixels within the same class, which is used in SFM as visualized in Fig. 3(a). Typically, it assigns high scores when two pixels are close or distant but belong to the same class, indicating similar embeddings. However, after class-level attention, each pixel-level embedding integrates contributions from redundant classes. This leads to decreased similarity among pixel-level embeddings of the same class, hindering spatial-level attention from effectively capturing dependencies on distant pixels within the same class. In Fig. 2, we visualize the spatial-level attention maps of different classes, confirming our conclusion that redundancy causes spatial-level attention to focus more on local regions and compromises the ability to capture long-range dependencies.
III-C3 Impact on latency
Without using CRM, the computational complexity of class-level and spatial-level attention is $O(K^{2}CH_{c}W_{c})$ and $O(KCH_{c}^{2}W_{c}^{2})$, respectively. With the help of CRM, the corresponding complexities become $O(P^{2}CH_{c}W_{c})$ and $O(PCH_{c}^{2}W_{c}^{2})$, which means that CRM can reduce the computation of approximately $1-\left({P}/{K}\right)^{2}$ and $1-{P}/{K}$ for class-level and spatial-level attention, respectively.
@@ III-C3 Impact on latency
TABLE II: Comparison with state-of-the-art methods. Img. and Txt. indicate whether the VLM’s image and text encoder are fine-tuned, respectively. A.B. signifies whether an additional backbone is employed.
@@
ZegFormer [11] [CVPR’22] ViT-B/16 COCO-Stuff 5.6 10.4 18.0 45.5 89.5
SimBaseline [17] [ECCV’22] ViT-B/16 COCO-Stuff 7.0 - 20.5 47.7 88.4
OVSeg [12] [CVPR’23] ViT-B/16 COCO-Stuff 7.1 11.0 24.8 53.3 92.6
DeOP [13] [ICCV’23] ViT-B/16 COCO-Stuff-156 7.1 9.4 22.9 48.8 91.7
SAN [14] [CVPR’23] ViT-B/16 COCO-Stuff 10.1 12.6 27.5 53.8 94.0
EBSeg [21] [CVPR’24] ViT-B/16 COCO-Stuff 11.1 17.3 30.0 56.7 94.6
SED [15] [CVPR’24] ConvNeXt-B COCO-Stuff 11.4 18.6 31.6 57.3 94.4
CAT-Seg [16] [CVPR’24] ViT-B/16 COCO-Stuff 12.0 19.0 31.8 57.5 94.6
ERR-Seg [ours] ConvNeXt-B COCO-Stuff 11.7 18.9 32.9 56.6 94.9
ERR-Seg [ours] ConvNeXt-B COCO-Stuff 13.9 20.5 35.8 56.9 95.4
OVSeg [12] [CVPR’23] ViT-L/14 COCO-Stuff 9.0 12.4 29.6 55.7 94.5
ODISE [19] [CVPR’23] ViT-L/14 COCO Panoptic 11.1 14.5 29.9 57.3 -
SAN [14] [CVPR’23] ViT-L/14 COCO-Stuff 12.4 15.7 32.1 57.7 94.6
FC-CLIP [20] [NeurIPS’23] ConvNeXt-L COCO Panoptic 14.8 18.2 34.1 58.4 95.4
SAM-CP [23] [ArXiv’24] ConvNext-L COCO-Panoptic - - 31.8 - -
Unpair-Seg [51] [ArXiv’24] ConvNext-L Merged-130K 14.6 19.5 32.6 52.2 93.0
EBSeg [21] [CVPR’24] ViT-L/14 COCO-Stuff 13.7 21.0 32.8 60.2 96.4
SED [15] [CVPR’24] ConvNeXt-L COCO-Stuff 13.9 22.6 35.2 60.6 96.1
CAT-Seg [16] [CVPR’24] ViT-L/14 COCO-Stuff 16.0 23.8 37.9 63.3 97.0
ERR-Seg [ours] ConvNeXt-L COCO-Stuff 15.8 22.7 36.5 60.7 96.6
ERR-Seg [ours] ConvNeXt-L COCO-Stuff 16.9 23.9 38.3 60.7 96.6
TABLE III: Efficiency comparison. The latency (ms) is computed on a single NVIDIA RTX 3090 GPU.

### v2

@@ 3.1.2 Redundant Class Reduction
The pipeline of the function $\Phi$ is illustrated in Fig. 4. $S$ represents the probability of each category at each spatial location. Subsequently, for each spatial location, the top $k$ classes are selected based on their probability values and arranged in descending order:
$N=Top(S,k),$ (4)
@@
where $N\in\mathbb{R}^{k\times H_{c}\times W_{c}}$ represents the indices of the top $k$ categories, $H_{c}\times W_{c}$ is the size of input cost maps. Then we calculate the scores for each category. The score for $t$-th category, denoted as $R(t)$, is computed as:
$R(t)=\sum_{c=1}^{k}{\lambda^{c-1}\left(\sum_{i=1}^{H_{c}}{\sum_{j=1}^{W_{c}}{\mathbb{I}\left[N(c,i,j)=t\right]}}\right)},$ (5)
@@
where $\lambda^{c-1}$ represents the score assigned to the class ranked $c$. $\lambda$ is the score coefficient and is expected to meet the condition $0<\lambda<1$, as higher-ranked categories should receive higher scores. The notation $\mathbb{I}[\cdot]$ represents a logical expression that yields 1 when the condition is true and 0 otherwise. The cost map used in $\Phi$ is $S_{4}$, since it is built from the origin final output of CLIP’s image encoder. After scoring all categories, we obtain $\left\{R(i)\right\}_{i=1}^{K}$, only the top $P$ categories are retained, thus forming redundancy-reduced $S_{3}$ and $S_{4}$.
3.1.3 Cost Maps Embedding and Merging
After removing redundant class, we employ a convolution layer with input channel 1 and output channel $C$ to embed the cost maps $S_{3}$ and $S_{4}$ into the latent space. To aggregate the hierarchical cost maps, we employ a feature pyramid strategy. Specifically, it can be computed as:
$S\leftarrow\textrm{conv}\left(S_{3}\right)+\textrm{up}\left(\textrm{conv}\left(S_{4}\right)\right),$ (6)
where $\mathrm{conv}$ denotes a convolution layer, $\mathrm{up}$ represents upsampling interpolation to ensure the same size. Subsequently add them to yield final cost maps $S\in\mathbb{R}^{P\times C\times H_{s}\times W_{s}}$, where $H_{s}=\frac{H}{16}$ and $W_{s}=\frac{W}{16}$.
3.2 Redundancy-Reduced Cost Aggregation
The cost-based OVSS framework employs spatial-level and class-level decoupled aggregation to enhance contextual information in the coarse-grained cost maps $S$. Although this aggregation is necessary and effective, it incurs significant computational redundancy due to the attention mechanism modeled across extensive sequence dimensions. Therefore, we introduce Spatial-level Sequence Reduction (SSR) and Class-level Sequence Reduction (CSR) strategies before the aggregation stages. SSR and CSR selectively compress the sequence lengths involved in attention computation, significantly accelerating context fusion while maintaining the segmentation accuracy.
3.2.1 Spatial-level Aggregation
Spatial-level aggregation aims to model the relationships between different spatial locations within the same class. The architecture of the spatial-level aggregation block is presented in Fig. 3 (a). The cost maps $S$ is first reshaped into a spatial-level sequence $S_{\mathrm{spa}}\in\mathbb{R}^{P\times(H_{s}\times W_{s})\times C}$ and then adopt self-attention mechanism, where the sequence length is $H_{s}\times W_{s}$. Notably, since the class vocabulary is dynamic in OVSS, the channel dimension $C$ is inherently class-agnostic, thus self-attention operates independently on each class sequence. Although this approach can effectively model intra-class relations, it requires computing self-attention over all $H_{s}\times W_{s}$ pixels for every class, leading to substantial computational redundancy.
Recognizing the inherent redundancy in modeling dense interactions across all $H_{s}\times W_{s}$ pixels per class, we introduce Spatial-level Sequence Reduction (SSR) to reduce the computational complexity in spatial-level aggregation. Specifically, it reduces the sequence length of $S$ before reshaping into a spatial-level sequence, as follows:
$S^{r}=f_{\mathrm{proj}}\left(\mathrm{conv}_{s=r_{1}}(S)\right),$ (7)
where $S^{r}\in\mathbb{R}^{P\times(\frac{H_{s}}{r_{1}}\times\frac{W_{s}}{r_{1}})\times C}$, $\mathrm{conv}_{s=r_{1}}$ is a convolution layer with stride $r_{1}$ that performs downsampling on the cost maps while increasing the channel dimension, and $f_{\mathrm{proj}}$ represents a linear projection layer that restores the original channel dimension. Subsequently, in the self-attention module, the original sequence $S_{\mathrm{spa}}$ serves as query to preserve output dimensions, and the compressed sequence $S_{\mathrm{spa}}^{r}$ functions as key and value.
The output of the self-attention module $S_{\mathrm{spa}}$ then passes through a depth-wise convolution layer, followed by a two-layer MLP. Considering SSR compresses the spatial dimension for computational efficiency, it is necessary to enhance the feature representation capacity within the MLP. Therefore, we replace the first linear layer of the MLP with a star operation [ma2024rewrite], leveraging its ability to effectively map features into a higher-dimensional space for richer interactions. Specifically, the modified MLP layer is defined as:
$S_{\mathrm{spa}}\leftarrow f_{\mathrm{out}}\left(\sigma\left(f_{\mathrm{in}}^{1}(S_{\mathrm{spa}})\odot f_{\mathrm{in}}^{2}(S_{\mathrm{spa}})\right)\right),$ (8)
where, $\sigma$ denotes the activation function, $f_{\mathrm{in}}^{1}$ and $f_{\mathrm{in}}^{2}$ are linear projections mapping from $\mathbb{R}^{C}$ to $\mathbb{R}^{C^{\prime}}$, $\odot$ represents element-wise multiplication, $f_{\mathrm{out}}$ is the output projection mapping back to $\mathbb{R}^{C}$, and $C^{\prime}$ is the expanded intermediate dimension.
3.2.2 Class-level Aggregation
Class-level aggregation aims to capture dependencies among different classes within the same spatial location. The architecture of our proposed class-level aggregation block is depicted in Fig. 3 (b). It begins by reshaping the cost maps $S$ into a class-level sequence $S_{\mathrm{cls}}\in\mathbb{R}^{(H_{s}\times W_{s})\times P\times C}$, where $P$ is the sequence length for the subsequent self-attention operation. The goal of this self-attention mechanism is to model inter-class relationships, facilitating the identification of the most relevant category for each pixel based on cross-class interactions. However, performing this attention computation over all $H_{s}\times W_{s}$ pixels per spatial location is computationally intensive and potentially redundant.
Recognizing that neighboring pixels often share similar features and thus exhibit correlated class dependencies, we introduce the Class-level Sequence Reduction (CSR) strategy. CSR applies an average pooling operation with kernel size $r_{2}\times r_{2}$ and stride $r_{2}$ to spatially downsample the cost maps before reshaping. This reduces the spatial dimensions to $\frac{H_{s}}{r_{2}}\times\frac{W_{s}}{r_{2}}$. Consequently, the reshaped class-level sequence becomes $S_{\mathrm{cls}}\in\mathbb{R}^{(\frac{H_{s}}{r_{2}}\times\frac{W_{s}}{r_{2}})\times P\times C}$. Subsequently, the modified MLP in spatial-level aggregation is also employed for improving feature representation capabilities. Finally, the output features are upsampled to the original spatial resolution ($H_{s}\times W_{s}$) and added element-wise to the input cost map $S$. This compression offers dual benefits: (1) it aggregates information from adjacent pixels into a single representation, providing richer context for category identification at class-level aggregation, and (2) it drastically reduces the sequence length involved in the self-attention computation from $H_{s}W_{s}$ to $\frac{H_{s}W_{s}}{r_{2}^{2}}$, significantly improving computational efficiency.
3.3 Upsampling Decoder
After the cost aggregation, the cost maps $S$ incorporate more contextual information but remain at low resolution, lacking fine-grained edge details. To recover high-resolution semantic maps, we leverage middle-layer visual features $\mathcal{F}_{i}^{v}$ extracted from the CLIP image encoder to guide the upsampling of $S$, following existing practices [cho2024cat]. Initially, $\mathcal{F}_{i}^{v}$ undergoes a convolution layer and is then concatenated with the cost maps at each class. The concatenated features are then passed through a refinement block, comprising two convolution layers. Finally, high-resolution cost maps $S_{hres}\in\mathbb{R}^{P\times C_{h}\times\frac{H}{4}\times\frac{W}{4}}$ is obtained, where $C_{h}$ is the channel dimension of the cost maps, which gradually decreases in the upsampling decoder. Finally, a linear project head project $S_{hres}$ into the size of $P\times(1)\times\frac{H}{4}\times\frac{W}{4}$, to get the final logits.
3.4 Analysis on Redundancy Reduction
A comprehensive analysis of our proposed redundancy reduction is provided in this section, encompassing three aspects: the quantitative analysis of class-level aggregation, the qualitative features of spatial-level aggregation, and the computational complexity.
3.4.1 Impact on class-level aggregation
@@
The original cost maps $S$ comprise the existing and redundant classes. Let $\mathcal{Q}_{e}=\{S(i)\}_{i=1}^{M}$ collect the original cost maps corresponding to the existing classes, where $M$ is the number of existing classes in $I$. The cost maps corresponding to the redundant classes are represented by $\mathcal{Q}_{r}=\{S(i)\}_{i=M+1}^{K}$, and those corresponding to the redundant classes after partial reduction are denoted as $\mathcal{Q}_{p}=\{S(i)\}_{i=M+1}^{P}$. In the subsequent subsections, we will analyze the impact of $\mathcal{Q}_{r}$ and $\mathcal{Q}_{p}$ on class-level and spatial-level aggregation (the primary operations on the cost maps), as well as on latency.
The class-level attention output of the $q$-th class in $\mathcal{Q}_{e}$ before reduction is:
$O^{e,r}(q)=\sum_{k=1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)},$
@@
$O^{r}(q)=\sum_{k=M+1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)}.$
@@
$\Delta_{p}=\frac{O^{p}(q)}{O^{e,p}(q)}<\Delta_{r}=\frac{O^{r}(q)}{O^{e,r}(q)}.$ (11)
@@
The above inequality suggests that as the number of redundant classes increases, they contribute more to the final output. Thus, existing classes contribute less. As such, we can deduce that reducing redundancy facilitates the class-level aggregation to focus more on existing classes. The proof is detailed in the Appendix.
3.4.2 Impact on spatial-level aggregation
@@
Spatial-level aggregation aims to capture dependencies among different pixels within the same class, it typically assigns high scores when two pixels are close or distant but belong to the same class, indicating similar embeddings. However, after class-level aggregation, each pixel-level embedding integrates contributions from redundant classes. This leads to decreased similarity among pixel-level embeddings of the same class, hindering spatial-level attention from effectively capturing dependencies on distant pixels within the same class. In Fig. 2, we visualize the spatial-level attention maps of different classes, confirming our conclusion that redundancy causes spatial-level attention to focus more on local regions and compromises the ability to capture long-range dependencies.
3.4.3 Impact on computational complexity
The computational efficiency of ERR-Seg is achieved through two redundancy reduction strategies: redundant class dimension reduction from $K$ to $P$, combined with spatial-level and class-level sequence reduction through downsampling ratios $r_{1},r_{2}$. For spatial-level attention, dimensionality compression transforms the spatial sequence from $K\times H_{s}W_{s}\times C$ to $P\times\frac{H_{s}W_{s}}{r_{1}^{2}}\times C$, reducing complexity from $O(KCH_{s}^{2}W_{s}^{2})$ to $O(PCH_{s}^{2}W_{s}^{2}/r_{1}^{2})$. For class-level attention, dimensionality compression transforms complexity from $O(K^{2}CH_{s}W_{s})$ to $O(P^{2}C\cdot H_{s}W_{s}/r_{2}^{2})$. This yields computational savings of $1-P/(r_{1}^{2}K)$ for spatial-level aggregation and $1-(P/(r_{2}K))^{2}$ for class-level aggregation.
4 Experiments
4.1 Experimental Setup
4.1.1 Traning dataset
Following the existing OVSS methods [cho2024cat, xie2024sed], we train our model on COCO-Stuff [caesar2018coco]. It has 118,287 images for training with 171 different classes.
4.1.2 Evaluation dataset
  • PASCAL-Context [mottaghi2014role] contains 459 classes with 5,104 images for validation. We report the scores of the full classes (PC-459) and the scores of the 59 most frequent classes (PC-59).
  • ADE20K [zhou2019semantic] has 2,000 validation images with 847 classes. Similarly to PASCAL-Context, we report the results of 150 frequent classes (A-150) and full 847 classes (A-847).
  • PASCAL VOC [everingham2010pascal] includes 1,449 images for validation with 20 common object classes. We report the performance of the 20 classes named PAS-20.
4.1.3 Analysis of the evaluation dataset
Prior studies have typically overlooked the label-level similarity between the validation and training sets, while it is necessary for open vocabulary-related tasks. In this section, we extract text embeddings for each category and then calculate the cosine similarity between the categories in the evaluation and training datasets, resulting in a similarity matrix. Subsequently, we map each category in the evaluation dataset to a category with the highest similarity score in the training dataset. The results are presented in Table 2; we report the proportion of labels in the validation sets with a similarity score equal to $1.00$, greater than $0.95$, $0.90$, and $0.85$.
Table 2: Label-level similarity between evaluation and training datasets.
Dataset $1.00$ $\geq 0.95$ $\geq 0.90$
PAS-20 0.45 0.90 1.00
PC-59 0.30 0.85 0.95
A-150 0.19 0.41 0.57
PC-459 0.10 0.21 0.29
A-847 0.06 0.12 0.21
Figure 5: Visualization of segmentation results on ADE20K-150 and ADE20K-847. The results of previous state-of-the-art methods SAN [xie2024sed] and CAT-Seg [cho2024cat] are also included for comparison.
4.1.4 Implementation details
We adopt the Convolutional CLIP [cherti2023reproducible] from OpenCLIP [ilharco_gabriel_2021_5143773], where the visual encoder is ConvNeXt [liu2022convnet] architecture. Our code is implemented on the top of Detectron2 [wu2019detectron2]. The AdamW [loshchilov2017decoupled] optimizer is employed, with an initial learning rate of $2\times 10^{-4}$ and a weight decay of $1\times 10^{-4}$. The learning rate is set to $2\times 10^{-6}$ for fine-tuning the CLIP model. We set $\lambda=0.1,k=3$ for RRHC and $r_{1}=2,r_{2}=2$ for RRCA. Our model is trained on COCO-Stuff using cross-entropy loss for 8080K iterations, with a batch size of $4$ and a crop size of $640\times 640$. $4$ NVIDIA RTX 3090 GPUs were used for training.
@@ 4.1.4 Implementation details
Table 3: Comparison with state-of-the-art methods. Img. and Txt. indicate whether the VLM’s image and text encoders are fine-tuned, respectively. A.B. signifies whether an additional backbone is employed.
@@
ZegFormer [ding2022decoupling] ViT-B/16 COCO-Stuff 5.6 10.4 18.0 45.5 89.5
SimBaseline [xu2022simple] ViT-B/16 COCO-Stuff 7.0 - 20.5 47.7 88.4
OVSeg [liang2023open] ViT-B/16 COCO-Stuff 7.1 11.0 24.8 53.3 92.6
DeOP [han2023open] ViT-B/16 COCO-Stuff-156 7.1 9.4 22.9 48.8 91.7
SAN [xu2023side] ViT-B/16 COCO-Stuff 10.1 12.6 27.5 53.8 94.0
EBSeg [shan2024open] ViT-B/16 COCO-Stuff 11.1 17.3 30.0 56.7 94.6
SED [xie2024sed] ConvNeXt-B COCO-Stuff 11.4 18.6 31.6 57.3 94.4
SCAN [liu2024scan] ViT-B/16 COCO-Stuff 13.2 10.8 30.8 58.4 97.0
MROVSeg [zhu2024mrovseg] ViT-B/16 COCO-Stuff 12.1 19.6 32.0 58.5 95.5
CAT-Seg [cho2024cat] ViT-B/16 COCO-Stuff 12.0 19.0 31.8 57.5 94.6
ERR-Seg (ours) ConvNeXt-B COCO-Stuff 11.7 18.9 32.9 56.6 94.9
ERR-Seg (ours) ConvNeXt-B COCO-Stuff 13.9 20.5 35.8 56.9 95.4
OVSeg [liang2023open] ViT-L/14 COCO-Stuff 9.0 12.4 29.6 55.7 94.5
ODISE [xu2023open] ViT-L/14 COCO Panoptic 11.1 14.5 29.9 57.3 -
SAN [xu2023side] ViT-L/14 COCO-Stuff 12.4 15.7 32.1 57.7 94.6
FC-CLIP [yu2024convolutions] ConvNeXt-L COCO Panoptic 14.8 18.2 34.1 58.4 95.4
Unpair-Seg [wang2024open] ConvNext-L Merged-130K 14.6 19.5 32.6 52.2 93.0
EBSeg [shan2024open] ViT-L/14 COCO-Stuff 13.7 21.0 32.8 60.2 96.4
SED [xie2024sed] ConvNeXt-L COCO-Stuff 13.9 22.6 35.2 60.6 96.1
MAFT+ [jiao2024maftp] ConvNeXt-L COCO-Stuff 15.1 21.6 36.1 59.4 96.5
OVSNet [liu2025stepping] ConvNeXt-L COCO-Panoptic 16.2 23.5 37.1 62.0 96.9
CAT-Seg [cho2024cat] ViT-L/14 COCO-Stuff 16.0 23.8 37.9 63.3 97.0
ERR-Seg (ours) ConvNeXt-L COCO-Stuff 15.8 22.7 36.5 60.7 96.6
ERR-Seg (ours) ConvNeXt-L COCO-Stuff 16.9 23.9 38.3 60.7 96.6
Table 4: Efficiency comparison. The latency (ms) is computed on a single NVIDIA RTX 3090 GPU.